# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v12

@@ D.7 Sensitivity to the Regional Heterogeneity in Endowment of Male Miners
I assess the influence of regional heterogeneity in female labor patterns. Nishinarita (1985) argues that, due to its geographical characteristics, coal mines in Hokkaidō (a northernmost island) tended to employ men from rural villages in the Tōhoku (northeastern) region of the main island as their main miners (see also Ogino 1993, p. 32). This forced the coal mines in Hokkaidō to use a sloping-face haulage method instead of using ato-yama. As a result, female miners in the Hokkaido mines were mainly engaged in the coal selection process rather than working in pits. It is unlikely that this regional heterogeneity could affect the location of mines, because the location of coal mines depends on the presence or absence of coal-bearing strata (Section 5.1). In addition, the number of mines in Hokkaidō is relatively small, meaning that such regional heterogeneity in labor supply is less likely to disturb the estimates. Despite this, to test whether such features lead to omitted variable bias, I added an indicator variable for the municipalities of the Hokkaidō prefecture in the regressions. Table D.3 shows the results from the specification, including an indicator variable for the municipalities in the Hokkaidō prefecture. The results are materially similar to those of the main results presented in the main text.
@@ D.8 Alternative Identification Strategy
I used IV estimation to obtain the baseline estimates. For the outcome variables that can be used as a panel dataset, an alternative estimation strategy may be the DID approach, which uses both cross-sectional and within variations for the identification. To implement the DID technique, municipalities that had been treated throughout the measured years would have to be excluded, while municipalities that had newer mines during the sampled periods were left. However, the number of municipalities with coal mines increased by only 23 between 1931 and 1936 (Online Appendix A.4 presents the raw number of mines). Consequently, as explained in Section 4.1 in detail, the proportion of treated municipalities decreased from $15$% to $13$% (Panel A of Table C.3). This feature of the assignments makes it difficult to employ the DID approach as the main identification strategy for the coal mine sample because the information that can be used for the identification is insufficient. [^63] Similarly, the number of municipalities with GSC mines increased by only 55 mines from 57 to 112 between 1931 and 1936 (see Online Appendix A.4 for the number of mines). The proportion of treated municipalities remained almost unchanged, from $5.0$% to $5.3$% (Panel A of Table C.3).
Nevertheless, I provide a brief case study using the DID approach for the subsample, including Ogawa coal mine in Iwate prefecture and the Miyoshiyama coal mine in Akita prefecture. As indicated in Figure 2b, both mines were located in the northeastern (Tōhoku) region, where the coal mines did not exist in 1931. In this light, it is expected to provide relatively clean experimental conditions for the DID estimation, albeit the available sample size is limited.
@@
where $\eta$, $\lambda$, and $\omega$ are the municipal fixed effect, year fixed effects, and random error term, respectively. $\textit{D}_{i,t}$ is a treatment indicator variable that takes one for the treatment group in the post-treatment period and zero for otherwise. $\mathbf{x}$ includes the accessibility measures on the stations and sea ports. Accessibility measure on the rivers is not included as it is a time-constant variable. However, the municipal fixed effect effectively controls for all these time-constant unobservable factors, including the geological strata and terrain that were potentially correlated with the location of mines.
Notice that this is a two-by-two DID setting for the TWFE model, which provides the unbiased DID estimator ($\hat{\delta}$) for the average treatment effect in the treated municipalities in the post-treatment period Abadie (2005). [^64] For a comprehensive review on the recent discussions on the TWFE-DID models, see de Chaisemartin and D’Haultfœuille (2022). The identification assumptions are the independence of the treatment, which is confirmed already, and the parallel trends assumption. Figure D.4 shows the time-series plots of the outcome variables used in the DID analysis for the Iwate and Akita subsample in Table D.4. Although there are moderate fluctuations due to the small sample size in each group, these outcomes do not show peculiar trends but show the materially similar trends in the pretreatment years (i.e., 1925 and 1930). The outcomes of the treatment group in the post-treatment period (i.e., 1935) show clear kinks in most cases, whereas those for the control group show marginal changes in the post-treatment period. A similar check for the mortality rates is unavailable because the data on deaths in 1925 are unavailable.
@@
Notes: This figure shows the time-series plots of the outcomes of the treatment and control groups in Iwate and Akita prefectures used in the regressions shown in Table D.4. Figures D.4aD.4bD.4cD.4e, and D.4d show the average number of people, average crude marriage rate, average marital fertility rate, average sex ratio and average household size, respectively. The data on the IMR in 1925 are unavailable.
@@
Notes: This table shows the results for the coal mines subsample in Iwate and Akita prefectures. Panel A: Column 1 shows the result for the log-transformed population. Columns 2–5 show the results for the crude marriage rate, marital fertility rate, male mortality rate, and female mortality rate, respectively.
Panel B: Columns 1-3 show the results for the infant mortality, fetal death, and child mortality rates, respectively. Columns 4–5 show the results for the sex ratio and average household size, respectively.
Table D.4 provides the results. In Panel A, column 1 shows the result for the local population, whereas columns 2–5 show those for the marital fertility rate and mortality rates, respectively. In Panel B, columns 1-3 present the results for the early-life mortality rates, whereas columns 4–5 show the results for the sex ratio and household size, respectively. Overall, the DID results align with the findings from my baseline results. First, the opening of these two coal mines increased the local population by approximately $8$%. Although the estimated magnitude is much smaller than the baseline magnitude listed in Table 2656565Notice that I compare my estimates with those for the results in 1935 or 1938 because the DID estimator captures the average treatment effect in the treated municipalities at the post-treatment period., this is rather consistent with the fact that both Ogawa and Miyoshiyama coal mines are considerably small mines with roughly 100 miners. [^66] In fact, this magnitude is rather similar to those estimated for the smaller coal mines in my expanded regressions (Online Appendix Table D.6). The estimates for fertility and mortality rates are also smaller than the baseline estimates, albeit the fertility pathway is still observed in the statistical sense. This aligns with the small influence of the small-scale mines (Online Appendix D.11). Moreover, both coal mines used few female miners: only 9 and 44 miners in Ogawa and Miyoshiyama coal mines, respectively. Thus, the regulations were less likely to influence these mines. To summarize, the local population growth in both mining areas was mainly led by the influx of male workers, which is supported by the estimate for the sex ratio and average household size (columns 4–5).
@@ D.9 Alternative Minerals: Heavy Metal Mines
I assess the validity of my main results on coal mines by comparing the results for heavy metal mines. Specifically, I focus on three heavy metals: gold, silver, and copper (GSC). This is because GSC mines comprise more than $80$% of all metal mines and generally extract these metals in one place. Although the AHC documented other types of metal mines, they were much smaller subsets: the number of lead and zinc, pig iron and steel, and tin mines recorded were only one, five, and four in 1931, respectively.
Historical evidence indicates that coal mining requires many more workers than heavy metal mining because coal is a bulky mineral (Section 2). Table C.3 indicates that the scale of GSC mines was much smaller than that of coal mines. In addition, unlike coal mines, the GSC mines were scattered across the archipelago and were less likely to agglomerate than coal mines (Online Appendix Figure C.2). The “enclave effect” suggests that some GSC mines are less likely to interact with local economies (Aragón and Rud 2013, 2015). [^67] A potentially relevant study is Stijns (2005); using a country-level dataset, the author provided suggestive evidence that the impact of resource abundance on socioeconomic outcomes may vary among resource types. This means that the impact of GSC mines on the regional population growth must be smaller than that of coal mines. To assess this, I compare the estimates of the reduced-form equation 2 for the coal and GSC mines because, unlike coal and fuel minerals, no specific stratum can predict the spatial distribution of GSC mines.
Table D.5 lists the results. Panel A summarizes the results for the main outcome variables. The estimates for the local population listed in columns 1 and 2 are $0.174$ for 1930 sample and $0.248$ for 1935 sample, respectively. These are less than half of the estimates for the coal mine sample (column 3 of Panels A and B in Table 2). This result confirms my baseline findings on the coal mines.
@@ D.11 Heterogeneity with Respect to Scale of Mines
While most of the municipalities with mines have a similar number of mines (i.e., one or two mines), several municipalities have many mines. [^68] In the 1930 (1935) sample, 85 (108) out of 93 (115) municipalities with mines have one to four mines, whereas the other 8 (7) municipalities have more than five mines. The average number of miners in the former group is roughly 1300 (1400) in 1930 (1935) sample, whereas that in the latter group is roughly 4000 (5400) in 1930 (1935) sample. Importantly, the causal effects are essentially heterogeneous across municipalities. This means that the municipalities with large-scale mine could have more significant impacts than those with small-scale mine. Estimating the heterogeneous treatment effects in detail is out of scope in this study. Estimating such varying effects also presents a practical issue: stratifying the treatment group disturbs the inference due to the small sample size in each subsample. Moreover, setting a certain threshold for the scale of mine is essentially difficult. Despite this, it may be useful to see whether the potential heterogeneity in the treatment effect concerning the scale of mines is reasonable. To assess the heterogeneity, I first stratify the treatment group into two subgroups using the median of the number of miners as a threshold. The key treatment indicator variable in equation 2 is then replaced with the two indicator variables for the two subgroups.
Table D.6 presents the results. Panels A and B show the results for the 1930 and 1935 samples, respectively. Columns 1–6 list the estimates for the local population, marital fertility, mortality rates, sex ratio, and average household size. Column 1 suggests that the relatively large-scale mines had more significant impacts on the local population growth. Fertility responses and mortality changes are clearly observed in the large-scale mines (columns 2 and 3). In addition to the female mortality, male mortality is improved in large-scale mines, which may reflect the wealth effects of the large mining area. The marginal effects for sex ratio and average household size are also greater for the large-scale mines.
@@
  • (28) Tuvia Blumenthal. Japanese economy in interwar period: International comparisons. In Takafusa Nakamura, editor, An Analysis of Japanese Economy in Interwar Period [in Japanese], pages 31–51. Tokyo, Tokyo, 1981.
@@
  • (30) Clément de Chaisemartin and Xavier D’Haultfœuille. Two-way fixed effects and differences-in-differences with heterogeneous treatment effects: a survey. Econometrics Journal, page utac017, June 2022.
@@
  • (75)
  • (76) Naoki Tanaka. Kindainihon tankōrōdōshi kenkyū (Research on the History of Coal Mining Labor in Modern Japan). [in Japanese]. Sōfuk̄an, Tokyo, 1984.
  • (77)

### v14

@@ D.7 Sensitivity to the Regional Heterogeneity in Endowment of Male Miners
I assess the influence of regional heterogeneity in female labor patterns. Nishinarita (1985) argues that, due to its geographical characteristics, coal mines in Hokkaidō (a northernmost island) tended to employ men from rural villages in the Tōhoku (northeastern) region of the main island as their main miners (see also Ogino 1993, p. 32). This forced the coal mines in Hokkaidō to use a sloping-face haulage method instead of using female miners (ato-yama). As a result, female miners in the Hokkaidō mines were mainly engaged in the coal selection process rather than working in pits. It is unlikely that this regional heterogeneity could affect the location of mines because the location of coal mines depends on the presence or absence of coal-bearing strata (Section 5.1). In addition, the number of mines in Hokkaidō is relatively small, meaning that such regional heterogeneity in labor supply is less likely to disturb the main estimates. Despite this, to test whether such features lead to omitted variable bias, an indicator variable for the municipalities of the Hokkaidō prefecture was added in the regressions. Table D.3 shows the results from the expanded regressions. The results are materially similar to those of the main results presented in the main text.
@@ D.8 Alternative Identification Strategy
The IV estimation was used to obtain the baseline estimates. For the outcome variables that can be used as a panel dataset, an alternative estimation strategy could be the DID approach, which uses both cross-sectional and within variations for the identification. To implement the DID technique, municipalities that had been treated throughout the measured years would have to be excluded. Simultaneously, municipalities that had newer mines during the sampled periods were left in. However, the number of municipalities with coal mines increased by only 23 between 1931 and 1936 (Online Appendix A.4 presents the raw number of mines). Consequently, as explained in Section 4.1 in detail, the proportion of treated municipalities decreased from $15$% to $13$% (Panel A of Table C.3). This feature of the assignments makes it difficult to employ the DID approach as the main identification strategy. This is because the information that can be used for the identification is insufficient. [^61] Similarly, the number of municipalities with GSC mines increased by only 55 mines from 57 to 112 between 1931 and 1936 (see Online Appendix A.4 for the number of mines). The proportion of treated municipalities remained almost unchanged, from $5.0$% to $5.3$% (Panel A of Table C.3).
Nevertheless, I provide a brief case study using the DID approach for the subsample, including Ogawa coal mine in Iwate prefecture and the Miyoshiyama coal mine in Akita prefecture. As indicated in Figure 2b, both mines were located in the northeastern (Tōhoku) region, where the coal mines did not exist in 1931. From this perspective, it is expected to provide relatively clean experimental conditions for the DID estimation, although the available sample size is limited.
@@
where $\eta$, $\lambda$, and $\omega$ are the municipal fixed effect, year fixed effect, and random error term, respectively. $\textit{D}_{i,t}$ is a treatment indicator variable that takes one for the treatment group in the post-treatment period and zero for otherwise. $\mathbf{x}$ includes the accessibility measures on the stations and sea ports. Accessibility measure on the rivers is not included as it is a time-constant variable. However, the municipal fixed effect effectively controls for all these time-constant unobservable factors, including the geological strata and terrain that were potentially correlated with the location of mines.
It should be noted that this is a two-by-two DID setting for the TWFE model. It provides the unbiased DID estimator ($\hat{\delta}$) for the average treatment effect in the treated municipalities in the post-treatment period (Abadie 2005). [^62] For a comprehensive review on the recent discussions on the TWFE-DID models, see de Chaisemartin and D’Haultfœuille (2022). The identification assumptions are the independence of the treatment, which has already been confirmed, and the parallel trends assumption. Figure D.4 shows the time-series plots of the outcome variables used in the DID analysis for the Iwate and Akita subsample in Table D.4. Although there are moderate fluctuations due to the small sample size in each group, these outcomes do not show peculiar trends but show materially similar trends in the pretreatment years (i.e., 1925 and 1930). The outcomes of the treatment group in the post-treatment period (i.e., 1935) show clear kinks in most cases, whereas those for the control group show marginal changes in the post-treatment period. A similar check for the mortality rates is unavailable because the data on deaths in 1925 are unavailable.
@@
Notes: This figure shows the time-series plots of the outcomes of the treatment and control groups in Iwate and Akita prefectures used in the regressions shown in Table D.4. Figures D.4aD.4bD.4cD.4e, and D.4d show the average number of people, average crude marriage rate, average marital fertility rate, average sex ratio and average household size, respectively. The data on the mortality in 1925 are unavailable.
@@
Notes: This table shows the results for the coal mines subsample in Iwate and Akita prefectures. Panel A: Column 1 shows the result for the log-transformed population. Columns 2–5 show the results for the crude marriage rate, marital fertility rate, male mortality rate, and female mortality rate, respectively. Panel B: Columns 1-3 show the results for the infant mortality, fetal death, and child mortality rates, respectively. Columns 4–5 show the results for the sex ratio and average household size, respectively.
Table D.4 provides the results. In Panel A, column 1 shows the result for the local population, whereas columns 2–5 show those for the marriage rate, marital fertility rate, and mortality rates. In Panel B, columns 1-3 present the results for the early-life mortality rates, whereas columns 4–5 show the results for the sex ratio and household size, respectively. Overall, the DID results align with the findings from the baseline results. First, the opening of these two coal mines increased the local population by approximately $8$%. Although the estimated magnitude is much smaller than the baseline magnitude listed in Table 2636363Notice that the estimates are compared with those for the results in 1935 or 1938 because the DID estimator captures the average treatment effect in the treated municipalities at the post-treatment period., this is consistent with the fact that both Ogawa and Miyoshiyama coal mines are small-scale mines with approximately $100$ miners. [^64] In fact, this magnitude is rather similar to those estimated for the smaller coal mines in my expanded regressions (Table D.6). The estimates for fertility and mortality rates are also smaller than the baseline estimates, albeit the fertility pathway is still observed in the statistical sense. This aligns with the small influence of the small-scale mines (Online Appendix D.11). Moreover, both coal mines used few female miners: only $9$ and $44$ miners in Ogawa and Miyoshiyama coal mines, respectively. Thus, the regulations were less likely to influence these mines. To summarize, the local population growth in both mining areas was mainly led by the influx of male workers, which is supported by the estimate for the sex ratio and average household size (columns 4–5).
@@ D.9 Alternative Minerals: Heavy Metal Mines
The validity of the main results on coal mines is assessed by comparing the results for heavy metal mines. Specifically, I focus on three heavy metals: gold, silver, and copper (GSC). This is because GSC mines comprise more than $80$% of all metal mines and generally extract these metals in one place. Although the AHC documented other types of metal mines, they were much smaller subsets: the number of lead and zinc, pig iron and steel, and tin mines recorded were only one, five, and four in 1931, respectively.
Historical evidence indicates that coal mining requires many more workers than heavy metal mining because coal is a bulky mineral (Section 2). Table C.3 indicates that the scale of GSC mines was much smaller than that of coal mines. In addition, unlike coal mines, the GSC mines were scattered across the archipelago and were less likely to agglomerate than coal mines (Figure C.2). The “enclave effect” suggests that some GSC mines are less likely to interact with local economies (Aragón and Rud 2013, 2015). [^65] A potentially relevant study is Stijns (2005); using a country-level dataset, the author provided evidence suggesting that the impact of resource abundance on socioeconomic outcomes may vary among resource types. This means that the impact of GSC mines on the regional population growth must be smaller than that of coal mines. To assess this, the estimates of the reduced-form equation 2 for the coal and GSC mines are compared because, unlike coal and fuel minerals, no specific stratum can predict the spatial distribution of GSC mines.
Table D.5 lists the results. Panel A (B) summarizes the results for the 1930 (1935) sample. The estimates for the local population listed in column 1 are $0.174$ for 1930 sample and $0.248$ for 1935 sample, respectively. These are less than half of the estimates for the coal mine sample (column 3 of Panels A and B in Table 2). This result confirms the baseline findings on the coal mines.
@@ D.11 Heterogeneity with Respect to Scale of Mines
While most of the municipalities with mines have a similar number of mines (i.e., one or two mines), several municipalities have many mines. [^66] In the 1930 (1935) sample, 85 (108) out of 93 (115) municipalities with mines have one to four mines, whereas the other 8 (7) municipalities have more than five mines. The average number of miners in the former group was approximately 1,300 (1,400) in 1930 (1935) sample, whereas that in the latter group was approximately 4,000 (5,400) in 1930 (1935) sample. Essentially, causal effects are heterogeneous across units. In my setting, this means that the municipalities with large-scale mine could have more significant impact than those with small-scale mines. Estimating the heterogeneous treatment effects in detail is beyond the scope of this study. Estimating such varying effects also presents a practical issue. Stratifying the treatment group disturbs the inference due to the small sample size in each subsample. Moreover, deciding a certain threshold for the scale of mine is essentially impossible. Despite this, it may be useful to determine whether the potential heterogeneity in the treatment effect concerning the scale of mines is reasonable. To assess the heterogeneity, the treatment group is first stratified into two subgroups using the median of the number of miners as a threshold. The key treatment indicator variable in equation 2 is then replaced with the two indicator variables for the two subgroups.
Table D.6 presents the results. Panels A and B show the results for the 1930 and 1935 samples, respectively. Columns 1–6 list the estimates for the local population, marital fertility, mortality rates, sex ratio, and average household size. Column 1 suggests that the relatively large-scale mines had more significant impacts on the local population’s growth. Fertility responses and mortality changes are clearly observed in the large-scale mines (columns 2 and 3). In addition to the female mortality, male mortality was improved in large-scale mines, which may reflect the wealth effects of the large mining area. The marginal effects for sex ratio and average household size are also greater for the large-scale mines.
@@
  • (6) Hunt Allcott and Daniel Keniston. Dutch disease or agglomeration? The local economic effects of natural resource booms in modern America. The Review of Economic Studies, 85(2):695–731, April 2018.
@@
  • (22) Dan Black, Terra McKinnish, and Seth Sanders. The economic impact of the coal boom and bust. The Economic Journal, 115(503): 449–476, April 2005.
@@
  • (24) Tuvia Blumenthal. Japanese economy in interwar period: International comparisons. In Takafusa Nakamura, editor, An Analysis of Japanese Economy in Interwar Period [in Japanese], pages 31–51. Tokyo, 1981.
  • (25)
  • (26) Francesco Caselli and Wilbur John Coleman II. The U.S. structural transformation and regional convergence: A reinterpretation. The Journal of Political Economy, 109(3): 584–616, June 2001.
  • (27)
  • (28) Clément de Chaisemartin and Xavier D’Haultfœuille. Two-way fixed effects and differences-in-differences with heterogeneous treatment effects: a survey. The Econometrics Journal, page utac017, June 2022.
@@
  • (47)
  • (48) Christian Siegel. Female relative wages, household specialization and fertility. Review of Economic Dynamics, 24: 152–174, 2017.
  • (49)
  • (51)
  • (53)
  • (54) Naoki Tanaka. Kindainihon tankōrōdōshi kenkyū (Research on the History of Coal Mining Labor in Modern Japan). [in Japanese]. Sōfūkan, Tokyo, 1984.
  • (55)
  • (56) James Tobin. Estimation of relationships for limited dependent variables. Econometrica, 26(1): 24–36, January 1958.
  • (57)
Documents used in the Appendices
  • (1)
  • (3)
  • (5)
  • (7)
  • (9)
  • (11)
  • (13)
@@
  • (24) Statistics Bureau of the Cabinet. Shōwagonen Kokuseichōsahōkoku, fukenhen (1930 Population Census of Japan, prefectural part). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1935a.
  • (25)